a canal boat travels 6 miles in 1 hr 30 mins. what is the average speed of the boat? give your answer in miles per hour.
Answer by mananth(16949)

1 hour 30 minutes = 6 miles
=1 1/2 hour
= 3/2 hours
..
r=d/t
r= (6 )/(3/2)
r= 6*2/3
r=4 mph
...
m.ananth@hotmail.ca